For the reaction CH4(g)  +2H2S(g) ⇌  CS2(g)  + 4H2(g)  , Kc = 0.046 at 687oC. In an experiment, 4.00 mol of CH4, 1.00 mol of H2S, 2.00mol of CS2, and 2.00 mol of H2 are mixed in a 500mL vessel. Is the reaction at equilibrium? If not, in which direction is it progressing?


1
Expert's answer
2022-02-22T14:26:03-0500

CH4(g)+2H2S(g)→CS2(g)+4H2(g).

Initial concentration.

CH4=1/0.25=4M

H2S=2/0.25=8M

CS2=1/0.25=4M

H2=2/0.25=8M


At equilibrium.

CH4=5.56M

Therefore the reaction is at equilibrium